A new game show requires a playing field with a perimeter of 54 yards and length 3 yards less than twice the width. What are the demensions?
Answer by checkley77(12844)

L=2W-3
2L+2W=54
2(2W-3)+2W=54
4W-6+2W=54
6W=54+6
6W=60
W=60/6
W=10 ANS. FOR THW WIDTH.
L=2*10-3=20-3=17 ANS. FOR THE LENGTH.
PROOF:
2*17+2*10=54
34+20=54
54=54